# Evaluation Guidelines for Empirical Studies in Software Engineering involving LLMs

### v1

@@ 4. Guidelines
Our guidelines focus on LLMs, that is, foundational models that use text as in- and output. We do not address multi-modal foundation models that support other media such as images. However, many of our recommendations apply to multi-modal models as well.
@@
The wording of our guidelines (must, should, may) follows RFC 2119 (Network Working Group, 1997) and RFC 8174 (Network Working Group, 2017). Throughout the following sections, we mention the information we expect researchers to report in the paper and/or in the supplementary material. We are aware that different publication venues have different page limits and that not all aspects can be reported in the paper. If information must be reported in the paper, we explicitly mention this in the specific guidelines. Of course, it is better to report essential information in the supplementary material than not at all. The supplementary material should be published according to the ACM SIGSOFT Open Science Policies (Graziotin, 2024).
At the beginning of each section, we provide a brief tl;dr summary that lists the most important aspects of the corresponding guideline. In addition to our recommendations, we provide examples from the SE research community and beyond, as well as the advantages and potential challenges of following the respective guidelines. We conclude each guideline by linking it to the study types. The remainder of this section is structured as follows:
@@ 4.1. Declare LLM Usage and Role
tl;dr: Researchers must disclose any use of LLMs in empirical studies in their paper. They should report their purpose, automated tasks, and expected benefits.
@@ 4.1.1. Recommendations
When conducting any kind of empirical study involving LLMs, researchers must clearly declare that an LLM was used. This should be done in a suitable section of the paper, for example, in the introduction or research methods section. For authoring scientific articles, this transparency is, for example, required by the ACM Policy on Authorship: The use of generative AI tools and technologies to create content is permitted but must be fully disclosed in the Work (Association for Computing Machinery, 2023). Beyond generic authorship declarations and declarations that LLMs were used as part of the research process, researchers should report the exact purpose of using an LLM in a study, the tasks it was used to automate, and the expected benefits in the paper.
@@ 4.1.2. Example(s)
The ACM Policy on Authorship suggests to disclose the usage of Generative AI tools in the acknowledgments section of the paper, for example: “ChatGPT was utilized to generate sections of this Work, including text, tables, graphs, code, data, citations” (Association for Computing Machinery, 2023). Similarly, the acknowledgments section could also be used to disclose GenAI usage for other aspects of the research, if not explicitly described in other sections. The ACM policy further suggests: “If you are uncertain about the need to disclose the use of a particular tool, err on the side of caution, and include a disclosure in the acknowledgments section of the Work” (Association for Computing Machinery, 2023). For double-blind review, during which the acknowledgments section is usually hidden, researchers can add a temporary “AI Disclosure” section in the same place the acknowledgments section would appear. An example of an LLM disclosure beyond writing support can be found in a recent paper by Lubos et al. (Lubos et al., 2024), in which they write in the methodology section:
@@ 4.1.4. Challenges
In general, we expect following our recommendations to be straightforward. One (perceived) challenge might be authors’ reluctance to disclose LLM usage for valid use cases such as improving language because they fear that AI-generated text makes reviewers think that the authors’ work is less original. However, the ACM Policy on Authorship is very clear in that any use of GenAI tools to create content must be disclosed.
@@ 4.2. Report Model Version, Configuration, and Customizations
This guideline focuses on documenting the model-specific aspects of empirical studies involving LLMs. While Section Report Tool Architecture beyond Models addresses how LLMs are integrated into larger systems and tools, here we concentrate on the models themselves, their version, configuration parameters, and any direct customizations applied to the model (e.g., fine-tuning). Whether only this guideline or also the following one applies for a particular study depends on the exact study setup and tool architecture. In any way, as soon as an LLM is involved, the information outlined in this guideline is essential to enable reproducibility and replicability.
@@ 4.2.1. Recommendations
A common customization approach for existing LLMs is fine-tuning. If a model was fine-tuned, researchers must describe the fine-tuning goal (e.g., improving the performance for a specific task), the fine-tuning procedure (e.g., full fine-tuning vs. Low-Rank Adaptation (LoRA), selected hyperparameters, loss function, learning rate, batch size, etc.), and the fine-tuning dataset (e.g., data sources, the preprocessing pipeline, dataset size) in the paper. Researchers should either share the fine-tuning dataset as part of the supplementary material or explain in the paper why the data cannot be shared (e.g., because it contains confidential or personal data that could not be anonymized). The same applies to the fine-tuned model weights. Suitable benchmarks and metrics should be used to compare the base model with the fine-tuned model (see Section Report Suitable Baselines, Benchmarks, and Metrics).
@@ 4.4. Report Prompts, their Development, and Interaction Logs
Prompts are critical for any study involving LLMs. Depending on the task, prompts may include various types of content, including instructions, context, and input data, and output indicators. For SE tasks, the context can include source code, execution traces, error messages, and other forms of natural language content. The output can be unstructured text or a structured format such as JSON. Prompts significantly influence the quality of the output of a model, and understanding how exactly they were formatted and integrated into an LLM-based study is essential to ensure transparency, verifiability, and reproducibility.
@@ 4.4.5. Study Types
Researchers must report the complete prompts that were used, including all instructions, context, and input data. This applies across all study types, but specific focus areas vary. For LLMs as LLMs as Annotators, researchers must document any predefined coding guides or instructions included in prompts, as these influence how the model labels artifacts. For LLMs as LLMs as Judges, researchers must include the evaluation criteria, scales, and examples embedded in the prompt to ensure a consistent interpretation. For LLMs used in LLMs for Synthesis tasks (e.g., summarization, aggregation), researchers must document the sequence of prompts used to generate and refine outputs, including follow-ups and clarification queries. For LLMs as LLMs as Subjects (e.g., simulating human participants), researchers must report any role-playing instructions, constraints, or personas used to guide LLM behavior. For Benchmarking LLMs for Software Engineering Tasks studies using pre-defined prompts (e.g., HumanEval, CruxEval), researchers must specify the benchmark version and any modifications made to the prompts or evaluation setup. If prompt tuning, retrieval-augmented generation (RAG), or other methods were used to adapt prompts, researchers must disclose and justify those changes, and should make the relevant code publicly available.
@@ 4.5.1. Recommendations
Researchers should use established reference models to compare humans with LLMs. For example, the reference model of Schneider et al. (Schneider et al., 2025) provides researchers with an overview of the design considerations for studies comparing LLMs with humans. If studies involve the annotation of software artifacts, and the goal is to automate the annotation process using LLM, researchers should follow systematic approaches to decide whether and how human annotators can be replaced. For example, Ahmed et al. (Ahmed et al., 2024) suggest a method that involves using a jury of three LLMs with 3 to 4 few-shot examples rated by humans, where the model-to-model agreement on all samples is determined using Krippendorff’s Alpha. If the agreement is high (alpha $gt$ 0.5), a human rating can be replaced with an LLM-generated one. In cases of low model-to-model agreement (alpha $\leq$ 0.5), they then evaluate the prediction confidence of the model, selectively replacing annotations where the model confidence is high ($\geq$ 0.8).
@@ 4.5.2. Example(s)
Ahmed et al. (Ahmed et al., 2024) evaluated how human annotations can be replaced by LLM generated labels. In their study, they explicitly report the calculated agreement metrics between the models and between humans. For example, the wrote that “model-model agreement is high, for all criteria, especially for the three large models (GPT-4, Gemini, and Claude). Table I indicates that the mean Krippendorff’s $\alpha$ is 0.68-0.76. Second, we see that human-model and human-human agreements are in similar ranges, 0.24-0.40 and 0.21-0.48 for the first three categories.”
@@ 4.5.4. Challenges
Measurement through human validation can be challenging. Ensuring that the operationalization of a desired construct and the method of measuring it are appropriate requires a deep understanding of (1) the construct, (2) construct validity in general, and (3) systematic approaches for designing measurement instruments (Sjøberg and Bergersen, 2023). Human judgment is subjective, which can lead to variability between judgments due to differences in experience, expertise, interpretations, and biases among evaluators (McDonald et al., 2019). For example, Hicks et al. found that “the struggle with adapting to AI-assisted work is more common for racially minoritized developers. That group also rated the quality of the output of AI-assisted coding tools significantly lower than other groups” (Hicks et al., 2025). Such biases can be mitigated with a careful selection and combination of multiple judges, but cannot be completely controlled for. Recruiting participants as human validators will always incur additional resources compared to machine-generated measures. Researchers must weigh the cost and time investment incurred by the recruitment process against the potential benefits for the validity of their study results.
@@ 4.6.1. Recommendations
Empirical studies using LLMs in SE, especially those that target commercial tools or models, should incorporate an open LLM as a baseline and report established metrics for inter-model agreement (see Section Report Suitable Baselines, Benchmarks, and Metrics). We acknowledge that including an open LLM baseline might not always be possible, for example, if the study involves human participants, and letting them work on the tasks using two different models might not be feasible. Using an open model as a baseline is also not necessary if the use of the LLM is tangential to the study goal.
@@ 4.6.3. Advantages
Using a true open LLM as a baseline improves the reproducibility of scientific research by providing full access to model architectures, training data, and parameter settings (see Section Report Model Version, Configuration, and Customizations), thereby allowing independent reconstruction and verification of experimental results. Moreover, by adopting an open-source baseline, researchers can directly compare novel methods against established performance metrics without the variability introduced by proprietary systems (see also Section Report Suitable Baselines, Benchmarks, and Metrics). The transparent nature of these models allows for detailed inspection of data processing pipelines and decision-making routines, which is essential for identifying potential sources of bias and delineating model limitations (see Section Report Limitations and Mitigations). Furthermore, unlike closed-source alternatives, which can be withdrawn or altered without notice, open LLMs ensure long-term accessibility and stability, preserving critical resources for future studies. Finally, the licensing requirements associated with open-source implementations lower financial barriers, making advanced language models attainable for research groups operating under constrained budgets.
@@ 4.6.4. Challenges
Open-source LLMs face several notable challenges (see also Section Report Limitations and Mitigations). First, they often lag behind the most advanced proprietary systems in common benchmarks, making it difficult for researchers to demonstrate clear improvements when evaluating new methods using open LLMs alone (see Section Report Suitable Baselines, Benchmarks, and Metrics). Additionally, deploying and experimenting with these models typically requires substantial hardware resources, in particular high-performance GPUs, which may be beyond reach for many academic groups (see Section Report Tool Architecture beyond Models).
@@ 4.6.5. Study Types
When evaluating LLMs for New Software Engineering Tools, researchers should use an open LLM as a baseline whenever it is technically feasible; if integration proves too complex, they should report the initial benchmarking results of open models (see Section Report Suitable Baselines, Benchmarks, and Metrics). In formal benchmarking studies and controlled experiments (see Section Benchmarking LLMs for Software Engineering Tasks), an open LLM must be one of the models under evaluation. For observational studies in which using an open LLM is impossible, investigators should explicitly acknowledge its absence and discuss how this limitation might affect their conclusions (see Section Report Limitations and Mitigations). Finally, when using LLMs for Synthesis, where an LLM serves to explore qualitative data or contrast alternative interpretations, researchers may report results of an open LLM.
4.7. Report Suitable Baselines, Benchmarks, and Metrics
@@
tl;dr: Researchers must justify all benchmark and metric choices in the paper and should summarize benchmark structure, task types, and limitations. Where possible, traditional (non-LLM) baselines should be used for comparison. Researchers must explain why the selected metrics are suitable for the specific study. They should report established metrics to make study results comparable, but can report additional metrics that they consider appropriate. Due the inherent non-determinism of LLMs, experiments should be repeated; the result distribution should then be reported using descriptive statistics.
@@ 4.7.1. Recommendations
When selecting benchmarks, it is important to fully understand the benchmark tasks and the expected results because they determine what the benchmark actually assesses. Researchers must briefly summarize why they selected certain benchmarks in the paper, They should also summarize the structure and tasks of the selected benchmark(s), including the programming language(s) and descriptive statistics such as the number of contained tasks and test cases. Researchers should also discuss the limitations of the selected benchmark(s). For example, many benchmarks focus heavily on isolated Python functions. This assesses a very specific part of software development, which is certainly not representative of the full breadth of software engineering.
@@
To compare traditional and LLM-based approaches or different LLM-based tools, researchers should report established metrics whenever possible, as this allows secondary research. They can, of course, report additional metrics that they consider appropriate. In any way, researchers must argue why the selected metrics are suitable for the given task or study.
@@
From a measurement perspective, researchers should reflect on the theories, values, and measurement models on which the benchmarks and metrics they have selected for their study are based. For example, a large open dataset of software bugs is a collection of bugs according to a certain theory of what constitutes a bug and the values and perspective of the people who labeled the dataset. Reflecting on the context in which these labels were assigned and discussing whether and how the labels generalize to a new study context is crucial.

### v2

@@ 5. Guidelines
Our guidelines focus on LLMs, that is, foundational models that use text as in- and output. We do not address multi-modal foundation models that support other media such as images. However, many of our recommendations apply to these models as well. We consider the extension of our guidelines to explicitly address multi-modal foundational models an important direction for future work.
@@
The wording of our guidelines (must, should, may) follows RFC 2119 (Network Working Group, 1997) and RFC 8174 (Network Working Group, 2017). Throughout the following sections, we mention the information we expect researchers to report in the paper and/or in the supplementary material. We are aware that different publication venues have different page limits and that not all aspects can be reported in the paper. If information must be reported in the paper, we explicitly mention this in the specific guidelines. Of course, it is better to report essential information that we expect in the paper in the supplementary material than not at all. The supplementary material should be published according to the ACM SIGSOFT Open Science Policies (Graziotin, 2024).
At the beginning of each section, we provide a brief tl;dr summary that lists the most important aspects of the corresponding guideline. In addition to our recommendations, we provide examples from the SE research community and beyond, as well as the advantages and potential challenges of following the respective guidelines. We conclude each guideline by linking it to the study types. The remainder of this section presents our eight guidelines and is structured as follows:
@@ 5.1. Declare LLM Usage and Role
tl;dr: Researchers must disclose any use of LLMs to support empirical studies in their paper. They should report their purpose, automated tasks, and expected benefits.
@@ 5.1.1. Recommendations
When conducting any kind of empirical study involving LLMs, researchers must clearly declare that an LLM was used (see Scope for what we consider relevant research support). This should be done in a suitable section of the paper, for example, in the introduction or research methods section. For authoring scientific articles, this transparency is, for example, required by the ACM Policy on Authorship: The use of generative AI tools and technologies to create content is permitted but must be fully disclosed in the Work (Association for Computing Machinery, 2023). Beyond generic authorship declarations and declarations that LLMs were used as part of the research process, researchers should report the exact purpose of using an LLM in a study, the tasks it was used to automate, and the expected benefits in the paper.
@@ 5.1.2. Example(s)
The ACM Policy on Authorship suggests to disclose the usage of GenAI tools in the acknowledgments section of the paper, for example: “ChatGPT was utilized to generate sections of this Work, including text, tables, graphs, code, data, citations” (Association for Computing Machinery, 2023). Similarly, the acknowledgments section could also be used to disclose GenAI usage for other aspects of the research, if not explicitly described in other sections. The ACM policy further suggests: “If you are uncertain about the need to disclose the use of a particular tool, err on the side of caution, and include a disclosure in the acknowledgments section of the Work” (Association for Computing Machinery, 2023). For double-blind review, during which the acknowledgments section is usually hidden, researchers can add a temporary “AI Disclosure” section in the same place the acknowledgments section would appear. An example of an LLM disclosure beyond writing support can be found in a recent paper by Lubos et al. (Lubos et al., 2024), in which they write in the methodology section:
@@ 5.1.4. Challenges
In general, we expect following our recommendations to be straightforward. One challenge might be authors’ reluctance to disclose LLM usage for valid use cases because they fear that AI-generated content makes reviewers think that the authors’ work is less original. However, the ACM Policy on Authorship is very clear in that any use of GenAI tools to create content must be disclosed. Before the introduction of LLMs, human proof-reading was possible and allowed and was not required to be declared. Our guidelines focus on research support beyond proof-reading and writing support (see Scope). However, over time, the threshold of what must be declared and what not will most likely evolve.
@@ 5.2. Report Model Version, Configuration, and Customizations
This guideline focuses on documenting the model-specific aspects of empirical studies involving LLMs. While Section Report Tool Architecture beyond Models addresses how LLMs are integrated into larger systems and tools, here we concentrate on the models themselves, their version, configuration parameters, and any direct customizations applied to the model (e.g., fine-tuning). Whether only this guideline or also the following one applies for a particular study depends on the exact study setup and tool architecture. In any event, as soon as an LLM is involved, the information outlined in this guideline is essential to enable reproducibility and replicability.
@@ 5.2.1. Recommendations
A common customization approach for existing LLMs is fine-tuning. If a model was fine-tuned, researchers must describe the fine-tuning goal (e.g., improving the performance for a specific task), the fine-tuning procedure (e.g., full fine-tuning vs. Low-Rank Adaptation (LoRA), selected hyperparameters, loss function, learning rate, batch size, etc.), and the fine-tuning dataset (e.g., data sources, the preprocessing pipeline, dataset size) in the paper. Researchers should either share the fine-tuning dataset as part of the supplementary material or explain in the paper why the data cannot be shared (e.g., because it contains confidential or personal data that could not be anonymized). The same applies to the fine-tuned model weights. Suitable benchmarks and metrics should be used to compare the base model with the fine-tuned model (see Section Use Suitable Baselines, Benchmarks, and Metrics).
@@ 5.4. Report Prompts, their Development, and Interaction Logs
Prompts are critical for any study involving LLMs. Depending on the task, prompts may include various types of content, including instructions, context, and input data, and output indicators. For SE tasks, the context can include source code, execution traces, error messages, and other forms of natural language content. The output can be unstructured text or a structured format such as JSON. Prompts significantly influence the quality of the output of a model, and understanding how exactly they were formatted and integrated into an LLM-based study is essential to ensure transparency, verifiability, and reproducibility. We remind the reader that these guidelines do not apply when LLMs are used solely for language polishing, paraphrasing, translation, tone or style adaptation, or layout edits (see Scope).
@@ 5.4.5. Study Types
Researchers must report the complete prompts that were used, including all instructions, context, and input data. This applies across all study types, but specific focus areas vary. For LLMs as Annotators, researchers must document any predefined coding guides or instructions included in prompts, as these influence how the model labels artifacts. For LLMs as Judges, researchers must include the evaluation criteria, scales, and examples embedded in the prompt to ensure a consistent interpretation. For LLMs for Synthesis tasks (e.g., summarization, aggregation), researchers must document the sequence of prompts used to generate and refine outputs, including follow-ups and clarification queries. For LLMs as Subjects (e.g., simulating human participants), researchers must report any role-playing instructions, constraints, or personas used to guide LLM behavior. For Benchmarking LLMs for Software Engineering Tasks studies using pre-defined prompts (e.g., HumanEval, CruxEval), researchers must specify the benchmark version and any modifications made to the prompts or evaluation setup. If prompt tuning, retrieval-augmented generation (RAG), or other methods were used to adapt prompts, researchers must disclose and justify those changes, and should make the relevant code publicly available.
@@ 5.5.1. Recommendations
Researchers should use established reference models to compare humans with LLMs. For example, the reference model of Schneider et al. (Schneider et al., 2025) provides researchers with an overview of the design considerations for studies comparing LLMs with humans. If studies involve the annotation of software artifacts, and the goal is to automate the annotation process using LLM, researchers should follow systematic approaches to decide whether and how human annotators can be replaced. For example, Ahmed et al. (Ahmed et al., 2025) suggest a method that involves using a jury of three LLMs with 3 to 4 few-shot examples rated by humans, where the model-to-model agreement on all samples is determined using Krippendorff’s Alpha. If the agreement is high (alpha $gt$ 0.5), a human rating can be replaced with an LLM-generated one. In cases of low model-to-model agreement (alpha $\leq$ 0.5), they then evaluate the prediction confidence of the model, selectively replacing annotations where the model confidence is high ($\geq$ 0.8).
@@ 5.5.2. Example(s)
Ahmed et al. (Ahmed et al., 2025) evaluated how human annotations can be replaced by LLM generated labels. In their study, they explicitly report the calculated agreement metrics between the models and between humans. For example, they wrote that “model-model agreement is high, for all criteria, especially for the three large models (GPT-4, Gemini, and Claude). Table I in their study indicates that the mean Krippendorff’s $\alpha$ is 0.68-0.76. Second, we see that human-model and human-human agreements are in similar ranges, 0.24-0.40 and 0.21-0.48 for the first three categories.”
@@ 5.5.4. Challenges
Measurement through human validation can be challenging. Ensuring that the operationalization of a desired construct and the method of measuring it are appropriate requires a deep understanding of (1) the construct, (2) construct validity in general, and (3) systematic approaches for designing measurement instruments (Sjøberg and Bergersen, 2023). Human judgment is subjective, which can lead to variability between judgments due to differences in experience, expertise, interpretations, and biases among evaluators (McDonald et al., 2019). For example, Hicks et al. found that “the struggle with adapting to AI-assisted work is more common for the racial minority developer cohort. That group also rated the quality of the output of AI-assisted coding tools significantly lower than other groups” (Hicks et al., 2025). Such biases can be mitigated with a careful selection and combination of multiple judges, but cannot be completely controlled for. Recruiting participants as human validators will always incur additional resources compared to machine-generated measures. Researchers must weigh the cost and time investment incurred by the recruitment process against the potential benefits for the validity of their study results.
@@ 5.6.1. Recommendations
Empirical studies using LLMs in SE, especially those that target commercial tools or models, should incorporate an open LLM as a baseline and report established metrics for inter-model agreement (see Section Use Suitable Baselines, Benchmarks, and Metrics). We acknowledge that including an open LLM baseline might not always be possible, for example, if the study involves human participants, and letting them work on the tasks using two different models might not be feasible. Using an open model as a baseline is also not necessary if the use of the LLM is tangential to the study goal.
@@ 5.6.3. Advantages
Using a true open LLM as a baseline improves the reproducibility of scientific research by providing full access to model architectures, training data, and parameter settings (see Section Report Model Version, Configuration, and Customizations), thereby allowing independent reconstruction and verification of experimental results. Moreover, by adopting an open-source baseline, researchers can directly compare novel methods against established performance metrics without the variability introduced by proprietary systems (see also Section Use Suitable Baselines, Benchmarks, and Metrics). The transparent nature of these models allows for detailed inspection of data processing pipelines and decision-making routines, which is essential for identifying potential sources of bias and delineating model limitations (see Section Report Limitations and Mitigations). Furthermore, unlike closed-source alternatives, which can be withdrawn or altered without notice, open LLMs ensure long-term accessibility and stability, preserving critical resources for future studies. Finally, the licensing requirements associated with open-source implementations lower financial barriers, making advanced language models attainable for research groups operating under constrained budgets.
@@ 5.6.4. Challenges
Open-source LLMs face several notable challenges (see also Section Report Limitations and Mitigations). First, they often lag behind the most advanced proprietary systems in common benchmarks, making it difficult for researchers to demonstrate clear improvements when evaluating new methods using open LLMs alone (see Section Use Suitable Baselines, Benchmarks, and Metrics). Additionally, deploying and experimenting with these models typically requires substantial hardware resources, in particular high-performance GPUs, which may be beyond reach for many academic groups (see Section Report Tool Architecture beyond Models).
@@ 5.6.5. Study Types
When evaluating LLMs for New Software Engineering Tools, researchers should use an open LLM as a baseline whenever it is technically feasible; if integration proves too complex, they should report the initial benchmarking results of open models (see Section Use Suitable Baselines, Benchmarks, and Metrics). In formal benchmarking studies and controlled experiments (see Section Benchmarking LLMs for Software Engineering Tasks), an open LLM must be one of the models under evaluation. For observational studies in which using an open LLM is impossible, investigators should explicitly acknowledge its absence and discuss how this limitation might affect their conclusions (see Section Report Limitations and Mitigations). Finally, when using LLMs for Synthesis, where an LLM serves to explore qualitative data or contrast alternative interpretations, researchers may report results of an open LLM.
5.7. Use Suitable Baselines, Benchmarks, and Metrics
@@
tl;dr: Researchers must justify all benchmark and metric choices in the paper and should summarize benchmark structure, task types, and limitations. Where possible, traditional (non-LLM) baselines should be used for comparison. Researchers must explain why the selected metrics are suitable for the specific study. They should report established metrics to make study results comparable, but can report additional metrics that they consider appropriate. Due to the inherent non-determinism of LLMs, experiments should be repeated; the result distribution should then be reported using descriptive statistics.
@@ 5.7.1. Recommendations
When selecting benchmarks, it is important to fully understand the benchmark tasks and the expected results because they determine what the benchmark actually assesses. Researchers must briefly summarize why they selected certain benchmarks in the paper, They should also summarize the structure and tasks of the selected benchmark(s), including the programming language(s) and descriptive statistics such as the number of contained tasks and test cases. Researchers should also discuss the limitations of the selected benchmark(s). For example, many benchmarks focus heavily on isolated Python functions. This assesses a very specific part of software development, which is certainly not representative of the full breadth of software engineering work (Chandra, 2025).
@@
To compare traditional and LLM-based approaches or different LLM-based tools, researchers should report established metrics whenever possible, as this allows secondary research. They can, of course, report additional metrics that they consider appropriate. In any event, researchers must argue why the selected metrics are suitable for the given task or study.
@@
From a measurement perspective, researchers should reflect on the theories, values, and measurement models on which the benchmarks and metrics they have selected for their study are based. For example, a large open dataset of software bugs is a collection of bugs according to a certain theory of what constitutes a bug, and the values and perspective of the people who labeled the dataset. Reflecting on the context in which these labels were assigned and discussing whether and how the labels generalize to a new study context is crucial.